# A High-Performance Evolutionary Multiobjective Community Detection Algorithm

### v1

@@ 4.1.2 Algorithms and Parameter Configuration
We compared HP-MOCD to two multi-objective evolutionary algorithms (MOEAs) and two classical greedy heuristics:
@@
  • MOCD [7]: Optimizes intra-community density and inter-community sparsity using the PESA-II framework.
  • MOGA-Net [13]: A modularity-based evolutionary algorithm with population-based search.
  • Louvain [6]: Greedy modularity maximization, widely used for its speed and simplicity.
@@
  • Leiden [9]: Improves Louvain by ensuring connected communities and more stable refinement.
All algorithms were configured based on their original publications. HP-MOCD used the following configuration, tuned empirically based on preliminary trials (see Section 3.2.3): crossover probability $C_{P}=0.8$, mutation probability $M_{P}=0.2$, and number of parents per crossover $E_{S}=4$. Table 1 summarizes the configurations.
@@
Table 1: Parameter settings and types of the compared algorithms. $C_{P}$ and $M_{P}$ denote crossover and mutation probabilities, and $E_{S}$ is the number of parents per crossover.
Algorithm Type Generations Population $C_{P}$ $M_{P}$ $E_{S}$ Reference
MOCD MOEA 100 100 0.9 0.1 - [7]
MOGA-Net MOEA 100 100 0.9 0.1 - [13]
Louvain Greedy - - - - - [6]
Leiden Greedy - - - - - [9]
@@
All algorithms were implemented in Python, with the exception of HP-MOCD, which is implemented in Rust and exposed to Python via PyO3333https://pyo3.rs/v0.25.0/. Although HP-MOCD benefits from parallel execution, none of the evaluated algorithms exploit parallelism in their algorithmic logic (e.g., population-level operators), making runtime comparisons consistent.
@@ 4.1.3 Evaluation Metrics
We evaluate performance using three complementary metrics:
Normalized Mutual Information (NMI):
This metric quantifies the agreement between the detected community partition $C$ and the ground-truth partition $C^{\prime}$ [31, 32], defined as:
@@
Adjusted Mutual Information (AMI):
AMI corrects the mutual information for chance [32], offering greater robustness when the number or size of communities varies:
@@
where $\text{MI}(U,V)$ is the mutual information between partitions $U$ and $V$, $\mathbb{E}[\text{MI}]$ is its expected value under random labeling, and $H(\cdot)$ denotes entropy.
4.1.4 Execution Environment
All experiments were executed on a Linux server equipped with a single Intel Core i7-4790 CPU @ 3.60 GHz (4 physical cores, 2 threads per core, 8 threads total) and 16 GB of RAM. HP-MOCD employed multithreaded execution, while all other algorithms were run in single-threaded mode to ensure a fair comparison. Each reported result corresponds to the mean over 20 independent runs initialized with different random seeds.
4.2 Impact of Network Size on Performance
Figure 5 presents a comparative evaluation of the algorithms’ performance across synthetic networks of increasing size, ranging from 10,000 to 100,000 nodes, with a fixed mixing parameter of $\mu=0.3$. Each result corresponds to the average over 20 independent runs, each initialized with a different random seed to account for stochastic variability and improve robustness. For all experiments, 95% confidence intervals were also computed to capture variability between runs.
(a) NMI Value
(b) AMI Value
(c) Execution Time (s)
Figure 5: AMI, NMI, and runtime (seconds) comparison of algorithms on LFR benchmark networks of different node counts ($n$) with a fixed mixing parameter ($\mu=0.3$), averaged over 20 independent runs.
The results show a clear advantage for HP-MOCD in terms of both detection quality and computational scalability. In Figures 5(a) and (b), HP-MOCD consistently achieves high NMI and AMI scores across all tested network sizes. Its detection quality remains stable and close to optimal, with NMI values near 1.0 and AMI values exceeding 0.9, indicating a strong alignment with the ground truth community structure. This consistency highlights HP-MOCD’s robustness to increasing graph size and suggests that the algorithm does not degrade as the problem scales. In contrast, the other MOEA-based methods (MOGA-Net and MOCD) demonstrate considerably lower accuracy at larger scales. Both exhibit modest improvements in NMI as $n$ increases, but plateau at significantly lower levels, and their AMI scores remain markedly below those of HP-MOCD. This confirms that the optimization strategies and genetic operators employed in HP-MOCD offer a more effective exploration of the solution space.
Figure 5(c) reveals the effects of network size on runtime. Note that the y-axis is presented on a logarithmic scale to accommodate the wide disparity in execution times. While the greedy algorithms (Louvain and Leiden) maintain a sublinear scaling as expected, their detection quality remains consistently inferior to that of HP-MOCD. Among the MOEA-based approaches, HP-MOCD exhibits a much lower runtime—up to many orders of magnitude faster than MOCD and significantly faster than MOGA-Net. This improvement can be directly attributed to HP-MOCD’s parallel architecture and efficient implementation. Importantly, when the network size exceeds 40,000 nodes, MOCD becomes impractical due to its steep runtime increase, whereas HP-MOCD remains computationally viable. These findings emphasize the algorithm’s scalability and practical utility in real-world settings where large graphs are common. In summary, HP-MOCD achieves the best balance between scalability and detection accuracy among the evaluated methods.
4.3 Resilience to Community Mixing
Figure 6 presents the performance of the algorithms under increasing levels of structural noise, as controlled by the mixing parameter $\mu$ in the LFR benchmark networks. Only algorithms with satisfactory performance from the previous experiment were included in this phase. MOCD and MogaNET algorithms that performed poorly at $\mu=0.3$ were excluded, as their performance is expected to deteriorate further with higher noise and their prohibitively long execution times, which exceeded 25 hours for a single run, making them impractical for this analysis.
Higher values of $\mu$ correspond to a greater proportion of edges connecting nodes across different communities, thereby reducing the separability of community structure and making the detection task more difficult. All experiments were conducted on networks with 100,000 nodes, averaged over 20 independent runs with different seeds. Results include 95% confidence intervals.
(a) NMI Value
(b) AMI Value
(c) Execution Time (s)
Figure 6: AMI, NMI and runtime (s) of the algorithms under comparison on LFR networks with different values of the mixing parameter $\mu$ and 100,000 nodes, averaged over 20 independent runs.

### v2

@@ 5.1 Comparison Algorithms and Parameter Configuration
We compare HP-MOCD against a diverse set of classical heuristics and multi-objective evolutionary algorithms, described below:
@@
  • MOCD [7]: A two-objective evolutionary method (based on PESA-II) optimizing intra-community density and inter-community sparsity.
  • MOGA-Net [16]: A modularity-oriented genetic algorithm evolving partitions directly.
  • Louvain [6]: A multilevel greedy heuristic that locally optimizes modularity.
@@
  • Leiden [10]: A refinement of Louvain ensuring connected communities and improved convergence.
  • ASYN-LPA [33]: A fast, parameter-free method based on label propagation among neighbors.
  • KRM [13]: An NSGA-III variant optimizing Kernel-k-means, Ratio-cut, and Modularity.
  • CCM [13]: An NSGA-III variant targeting Community Fitness, Community Score, and Modularity.
  • CDRME [14]: A decomposition-based evolutionary algorithm using spectral-based seeding and optimizing multiple modularity-related objectives.
All algorithms were implemented in Python, except HP-MOCD, which was implemented in Rust and exposed to Python via PyO3222https://pyo3.rs/v0.25.0/. Although HP-MOCD exploits parallelism in population-level operators, runtime comparisons reported in Sections 5.2.1 and 5.2.2 were performed in single-threaded mode except for HP-MOCD. Single-thread and parallel performance are evaluated separately in Appendix B. The algorithm baseline configurations are summarized in Table 1, based on their original publications.
The evolutionary process in HP-MOCD is governed by two key hyperparameters: population size and number of generations. To determine appropriate values, we conducted empirical experiments on synthetic benchmarks, using the same configuration as in our main experiments, except for varying $n$ (see Section 5.2). We analyzed the convergence behavior of the algorithm under these settings. As shown in Figure 5, 100 generations provided a favorable trade-off between solution quality and computational cost; for the other parameters and scenarios, we leave them in the Appendix C. HP-MOCD used the following configuration: crossover probability $C_{P}=0.8$, mutation probability $M_{P}=0.2$, and number of parents per crossover $E_{S}=4$. Table 1 summarizes the configurations.
(a) $maxgen=5$
(b) $maxgen=50$
(c) $maxgen$ = 100
Figure 5: Performance results across varying $maxgen$ values with $\mu=0.1$. Refer to Appendix C for details on the experimental setup.
@@
Table 1: Parameter settings and types of the compared algorithms. “Gens" and “Pop" denote the number of generations and population size for evolutionary algorithms. $C_{P}$, $M_{P}$, and $E_{S}$ refer to crossover probability, mutation probability, and number of parents per crossover, respectively.
Algorithm Type Gens Pop $C_{P}$ $M_{P}$ $E_{S}$ Reference
ASYN-LPA Label Prop. [33]
Louvain Greedy [6]
Leiden Greedy [10]
MOGA-Net MOEA 100 100 0.9 0.1 [16]
MOCD MOEA 100 100 0.9 0.1 [7]
KRM MOEA 200 100 0.9 0.1 [13]
CCM MOEA 200 100 0.9 0.1 [13]
CDRME MOEA 100 100 0.9 0.2 [14]
@@ 5.1.1 Evaluation Metrics
We assess the similarity between the detected and ground-truth community structures using three widely adopted external validation metrics: Normalized Mutual Information (NMI), Adjusted Mutual Information (AMI), and the pairwise F1-score [34, 24]. These metrics evaluate how well the detected partition recovers the original grouping of nodes, from both an information-theoretic and structural perspective. Throughout our experiments, these metrics are used in a complementary fashion to capture different aspects of clustering quality.
Unlike standard classification tasks, community detection does not assume fixed label identities—only the grouping of nodes into communities matters. As such, traditional per-label precision and recall-based F1-scores are not applicable. Instead, we adopt a pairwise evaluation framework, in which all unordered node pairs are treated as binary decisions such that each pair is considered positive if both nodes belong to the same community, and negative otherwise. Precision and recall are computed by comparing these pairwise co-assignment decisions in the predicted and ground-truth partitions, and the F1-score is their harmonic mean [34]. Each metric captures a different aspect of clustering quality. We use them jointly to ensure a comprehensive assessment. Below, we describe each of these metrics in detail.
  • Normalized Mutual Information (NMI): This metric quantifies the agreement between the detected community partition $C$ and the ground-truth partition $C^{\prime}$ [35, 36], defined as:
@@
  • Adjusted Mutual Information (AMI): AMI corrects the mutual information for chance [36], offering greater robustness when the number or size of communities varies:
@@
    where $\text{MI}(U,V)$ is the mutual information between partitions $U$ and $V$, $\mathbb{E}[\text{MI}]$ is its expected value under random labeling, and $H(\cdot)$ denotes entropy. Like NMI, AMI ranges from 0 to 1 but is less sensitive to random or imbalanced partitions.
  • F1-Score: This metric evaluates clustering accuracy by analyzing all unordered node pairs. Each pair is treated as a binary classification, where it is labeled positive if both nodes belong to the same community, and negative otherwise. Comparing these assignments between the predicted and ground-truth partitions, we compute [37, 34]:
[…]
    where precision is the fraction of predicted intra-community pairs that are also present in the ground truth, and recall is the fraction of true intra-community pairs correctly recovered. This pairwise evaluation captures the global consistency of group assignments and is invariant to label permutations.
  • Modularity: Although modularity is not an external validation metric, it remains widely used to assess the internal structure of communities based on edge density [28]. It compares the observed number of edges within communities to the expected number in a randomized null model that preserves the degree distribution. In essence, modularity quantifies how much more densely connected the nodes are within communities than would be expected by chance. Its score theoretically ranges from $-\frac{1}{2}$ to 1, with higher values indicating stronger community structure. It is formally defined as:
    where $A_{ij}$ is the adjacency matrix, $k_{i}$ and $k_{j}$ are the degrees of nodes $i$ and $j$, $m$ is the total number of edges, and $\delta(c_{i},c_{j})$ is 1 if nodes $i$ and $j$ are assigned to the same community and 0 otherwise. We include modularity in our evaluation to allow direct comparison with modularity-optimizing methods such as Louvain and Leiden.
5.2 Synthetic Benchmark Networks
We use the LFR benchmark generator [38], as implemented in the networkx Python library [^3] https://networkx.org/documentation/stable/reference/generated/networkx.generators.community.LFR_benchmark_graph.html, to generate synthetic networks with known ground-truth communities. LFR is widely adopted due to its ability to model realistic topological characteristics such as power-law degree distributions and heterogeneous community sizes. We adopt two main configurations:
  • Robustness test: Fixing network size $n=1{,}000$, we vary the mixing parameter $\mu\in[0.1,0.8]$ in increments of 0.1. This parameter controls the fraction of links that connect each node to other communities, such that lower values imply well-separated clusters, while higher values introduce more structural ambiguity.
  • Scalability test: Network size $n\in\{500,1{,}000,1{,}500,2{,}000,\ldots,20{,}000\}$, with average degree 20 and maximum degree 50. Community sizes vary in [20,100][20,100][ 20 , 100 ], and other parameters are fixed as: degree exponent $\tau_{1}=2.5$, community size exponent $\tau_{2}=1.5$, no overlapping nodes, and one community per node.
All configurations were replicated with 20 different random seeds, and average scores with standard deviation are reported.
5.2.1 Robustness to Community Mixing
Figure 6 presents the performance of the algorithms under increasing levels of structural noise, controlled by the mixing parameter $\mu$ in LFR benchmark networks. This parameter determines the proportion of edges that connect nodes to different communities: lower $\mu$ values represent well-separated clusters, while higher values introduce greater inter-community connectivity, making the detection task more challenging.
All evaluations were conducted on synthetic networks with 1,000 nodes, using 20 independent seeds for each $\mu\in\{0.1,0.2,\ldots,0.8\}$. For each configuration, we report the mean values and 95% confidence intervals of the evaluation metrics.
(a) NMI
(b) AMI
(c) Modularity
(d) Execution Time (s)
Figure 6: Performance of the evaluated algorithms on LFR networks with varying levels of structural noise ($\mu$), using 1,000 nodes. Each result represents the mean over 20 runs. Metrics shown include: (a) Normalized Mutual Information (NMI), (b) Adjusted Mutual Information (AMI), (c) modularity, and (d) execution time in seconds. Shaded areas indicate 95% confidence intervals.